OpenStudy (anonymous):

Motion Problem: At 4:00 Gigi and Madel left the music house. Gigi walked east and reached her house at 5:00. Madel rode her bike west and reached her house at 4:15. Madel traveled twice as fast as Gigi. Their houses are 9km apart. How far didi each travel?

OpenStudy (anonymous):

okay, lets think about this. gigi walked for an hour to reach her house madel rode for 15 minutes and reached her house. their houses are 9 km apart

OpenStudy (anonymous):

lets say gig's speed was S km/hour

OpenStudy (anonymous):

we know that distance = speed x time

OpenStudy (anonymous):

ans is 3 km

OpenStudy (anonymous):

so gigi traveled S km in one hour

OpenStudy (anonymous):

if speed of didi is x... x+2x/4=9

OpenStudy (anonymous):

madel's speed is 2S. in 15 minutes, that is, a quarter hour, she travels 2S x 1/4 = S/2 km

OpenStudy (anonymous):

thanks :]

OpenStudy (anonymous):

total distance is 9 km so S+S/2 = 9km S = 6km that means that gigi's house is 6 km away from the music store and madel's house is 3 km away from the music store